# A Large-Scale Privacy Assessment of Android Third-Party SDKs

### v1

@@ 5.2. RQ2: Privacy Collection and Potential Leakages at the SDK level
As shown in Table 4, location and sensors’ data (category C3) reserves the greatest portion of collection actions (reflected in the total number of caught access traces in the Table), contributing 1002 out of 2318 (43.2%) caught collection actions. The remaining privacy data categories ordered by their number of caught collection actions are data for wireless communication (C2, 22.8%), data for chip, cellular and peripheral (C1, 17.9%), media and software-specific data (C4, 13.4%), and personal data (C5, 2.8%), respectively.
From the privacy data sharing perspective, the access of wireless communication data (C2) constitutes the largest part of the tainted flow, as 172 out of 346 (49.7%) traces are tainted. The numbers of sharing actions involving data in categories C3 and C1 are following up, reserving 37.6% and 12.7% of all sharing actions, respectively. No sharing actions about data in categories C4 and C5 are observed. From the identifiability perspective, we find that most privacy data involved in sharing actions are referred to as personally identifiable information (PII) in relevant literature (ren2018longitudinal, ; ren2016recon, ; shen2021understanding, ; wang2020exploring, ).
We remark that whether SDKs can collect privacy data eventually depends on the permission and privilege of their embedding apps, especially for the collection and sharing of PIIs. We also highlight that the tainted traces do not automatically imply the potential exfiltration in runtime, which we will discuss later in this section.
@@
Figure 3. A false positive example of SSID reading (line 7)
@@
During our analysis, we noticed that static taint analysis cannot well determine the execution paths in conditional and branched code. If any tainted trace is found spanning over an unreachable, dead, or logically infeasible code, we exclude the tainted trace from the statistics of potential leakages. In Figure 3, we provide an example code reverse-engineered from one of our examined SDKs named “mipush”. We find it attempts to read the SSID of the connected WiFi but is later assessed by us as a false alarm. The SSID reading action (can be realized by the method getExtraInfo (line 7) until Android 10) would only take place if the device is not connecting to WiFi (after line 4), therefore we determine this is a logically infeasible code regarding the privacy data access. In this study, we manually exclude the false alarms from the 346 traces and manage to identify 338 (97.7%) logically valid traces, among which each trace indicates an occurrence of privacy leakage issue.
For the sharing operations, we observe that 332 out of 338 (98.2%) valid traces are tainted at network interfaces. The six remaining traces are tainted at system setting APIs. Compared with sharing privacy data through the Internet, we remark that writing privacy data in system settings is a more worrying issue because it exploits the privilege of pre-installed apps and is capable of sharing user privacy even without reading any privacy data from the Android OS. We will detail this in Section 6.
5.3. RQ3: SDK-level Privacy Compliance
We resort to cross-checking the privacy policies with the actual data collection behaviors of the collected SDKs to understand the status quo of privacy compliance at the SDK level. We remark our assessment only focuses on the 109 out of 158 (69.0%) SDKs that provide privacy policies, and the 27 data types that appear in the data collection behaviors observed from the taint analysis. We present our assessment results in Table 5.
Table 5. Assessment of privacy compliance by data types
@@ 5.3. RQ3: SDK-level Privacy Compliance
Our evaluation finds that 40 out of 109 (36.7%) SDKs collected excessive privacy data compared with their privacy policies (e.g., an SDK collects personal identifiers but does not declare them in its privacy policy). Among the 40 SDKs that excessively collect users’ privacy, the number of undeclared data types ranges from 1 to 9. Each SDK has an average of three undeclared data types. As shown in Table 5, 22 out of the 41 pre-identified data types (81.5%) have been excessively collected by the examined SDKs. The app list, clipboard data, and location are the top three most excessively collected privacy data types. Specifically, 22 SDKs (20.2%) collect the app list without any declaration in their privacy policies, followed by clipboard data and location that are excessively collected by 11 (10.0%) SDKs. By jointly considering data collection behaviors (refer to Table 4), our findings show that the app list and clipboard data are the two most stealthily collected data types. Only 4 SDKs collecting the app list have mentioned it in their privacy policies (leading to 22 excessive collection behaviors). None of the 11 caught SDKs have declared to read the clipboard data.
@@
In addition, our study unveils two interesting facts. First, we find that some SDK developers tend to use vague or general terminologies to describe collected privacy data such as device identifiers, advertising identifiers, and device information instead of listing out specific data collected. Developers’ interpretation of these ambiguous terminologies greatly differs. For example, a few developers define “identifiers” as advertising identifiers although the majority use this terminology to represent device identifiers, resulting in over-claiming cases in device identifiers but excessive-claiming cases in Android ID and Google Ad ID. The second fact is that many developers fail to clearly specify the sources of collected data types, especially the location. Commonly, location can be obtained by accessing GPS in mobile devices or inferred from the information of the connected network and IP addresses. Our analysis of privacy policies finds that many SDKs tend to declare their access to both the GPS and IP addresses to determine locations. However, our taint analysis shows that most SDKs turn to invoke location APIs only, leading to many over-claiming cases of IP addresses. This phenomenon may confuse the end users in understanding which exact data types are collected from them, resulting in poor transparency of the privacy policies.
Responsible Disclosure
We have reported all our findings to the relevant parties and we also have kept them confidential for at least 90 days. For each SDK that found a compliance issue, we manually searched the contact left by its developers and e-mailed our findings since our first round of SDK investigation. Although we observed slight mitigation from re-inspection, we did not receive positive responses regarding non-compliance concerns from the developers that we have attempted to contact. We also expressed our concern to Google about the system settings injection issues (to be detailed in Section 6) and suggested removing the pre-installed privileges in writing system settings in the future release. However, we are told that the pre-installed privilege is not regarded as a vulnerability and the relevant permission mechanism would remain unchanged. It is at the device manufacturer’s discretion to select and grant pre-installed privileges to third-party apps.
6. System Settings Injection Issues
Figure 4. Information flow demonstrates how the device ID is accessed and saved to the system settings by one of our analyzed SDKs (the obfuscated code is obtained from reverse engineering and has been simplified to save space)
Table 6. List of system setting entries found in our investigation that contains a UUI
SDK name Entry name Description of the value Operations to System Settings
[…]
Baidu LBS com.baidu.deviceid.v2 App’s package name and Android ID, MD5 $\ddagger$ Read only
[…]
$\dagger$ depending on the permissions granted to the embedded app. $\ddagger$ inferred description based on the reading logic in the implementation.
During our analysis, we find some traces from an SDK named “Alicloud push” are sunk at writing a value into the system settings. Specifically, two entries of system settings are involved, as shown in Table 6. Writing into the system settings requires privileged permission so that ordinary third-party apps are unable to do so. However, we find that SDKs being embedded by pre-installed apps [^6] We provide more technical details of the pre-installed apps and permission for writing into system settings in Appendix A.4. can write arbitrary values into the system settings even without user’s awareness. We name this finding a system settings injection issue, because it can be manipulated by an SDK to create and share user-unresettable identifiers (UUIs).
Although leaking privacy data through covert channels has been studied in existing literature (reardon2019fifty, ), the system settings injection issues still seriously concern us as the caught SDK subtly evades accessing highly sensitive UUIs (e.g., IMEI, as discussed in  (reardon2019fifty, )) but attempts to generate UUIs on their own. The caught SDK then takes advantage of its popularity and abuses the “preinstalled” privilege by counting on itself to be embedded in pre-installed apps or even privileged system apps. By exploiting this issue, the system settings become a covert channel to enable privacy-sensitive data sharing to apps without privileged permission.
Considering the exploitability of the pre-installed privileges, we carry out another round of investigation of all 158 SDKs to find if there exists any other potential privacy infringement by taking advantage of the system settings mechanism. As a result, we managed to find another two entries that have been read in an SDK named “Baidu LBS” (see Table 6). As their names imply, these two entries are obviously written by third-party privileged apps. We suspect the formation of the two entries should follow a similar routine, and accordingly reverse engineer the involved SDK and present our inferred descriptions of the two entries in Table 6. Unfortunately, we did not find any app or pre-installed app to write these entries into the system settings. Who exploited this system settings issue remains unknown to us.
Impact
We notice that the two studied SDKs are widely adopted in third-party apps, especially in the Chinese market. Therefore, we conduct a small-scale study to explore the impact of the system settings injection issues. We collect devices from 9 OEM manufacturers (8 manufacturers’ devices purchased in the Chinese market and 1 Google Pixel device purchased in the US market, please refer to Appendix A.5 for the full list), download the packages of pre-install apps, and investigate whether the two studied SDKs have ever been embedded in any of the pre-installed apps. We observe a worrisome fact that except for the Google Pixel devices, all the tested devices have at least one pre-installed app embedding the investigated SDKs. We assume that the choice of pre-installed apps should remain the same across different device models produced by a manufacturer. By referring to the latest market report (statcounter2024worldwide), these 8 OEM manufacturers reserve over 66% global Android smartphone market share. They also shipped over 94% new Android devices in China in the year 2023 (statcounter2024china, ), implying that almost all Android smartphones sold in the Chinese market pose a privacy leakage risk.
7. Privacy Compliance Re-inspection
Our study aims to unveil the comprehensive landscape of privacy compliance at the SDK level. Instead of conducting a one-shot analysis, we carry out a re-inspection to explore the trend of privacy collection among the same group of the tested SDKs. More specifically, on the basis of our large-scale analysis performed on 158 SDKs collected in October 2022, we re-visit the two sources and collect the same group of SDKs in their latest release version as of October 2023. Thus, we conduct a cross-version analysis of the privacy collection and sharing practices of the examined SDKs.
@@ 7. Privacy Compliance Re-inspection
Table 7. Changes in the tainted traces observed from re-inspection after 12 months (only tainted data types are displayed)
@@
We assess the same group of SDKs in their latest release versions and summarize the tainted traces in Table 7. We managed to find 311 traces tainted from the latest version of 158 SDKs. Compared with our initial assessment, we observed 35 (10.1%) fewer traces that attempt to collect user privacy data and share them in public spaces.
@@
The change between the outcomes of the two assessments mainly falls in the sources of tainted traces, which correspond to what types of privacy data have been collected and shared. As shown in Table 7, we find obvious growth from the sharing of data relevant to media and installed software (category C4), reserving over 1/3 of the total caught traces. Among the 105 tainted traces, reading and sharing the list of installed apps ranks the top among all sources, evidenced by 53 traces related to this type of privacy data. Although the access of data in that category has been pervasively observed in our initial assessment (310 traces that are not tainted, as shown in Table 4), our taint analysis did not catch any traces of sharing such data.
Additionally, we find that the location and SSID/BSSID contribute to the two greatest declines in the number of collection behaviors. We also observe an overall decrease in the collection of UUIs, including device identifiers (category C1) and identifiers of hardware components for wireless communication (category C2). By jointly considering the historical changes in permission requirements, we find that the access rules of most data that show a decrease in collection behaviors have undergone a significant update in Android 10, in which the third-party apps’ access to personally identifiable data has been strictly restricted (google2023privacy10, ). Specifically, third-party apps’ access to UUIs is prohibited. Therefore, the implementation to collect UUIs becomes invalid unless they are embedded in system apps. Permission to access data that may leak users’ physical location, such as GPS location, SSID, and BSSID, has also been escalated. A finer location permission becomes mandatory. Such permission escalation makes SDKs unable to stealthily access the aforementioned data unless the embedding apps explicitly request permission from users.
We also observe that the trends in SDKs’ privacy compliance show hysteresis along with the evolution of Android OS, given Android 10 was released in 2019. Such hysteresis could also be reflected in the data type demonstrating the largest growth of tainted traces, namely the app list. We consider SDK developers to treat the app list as an alternative to the UUIs and share it for user identification and profiling purposes. However, we find the SDK developers’ efforts in collecting user privacy lag behind Google’s strengthening privacy protection mechanism of Android OS, since access to the app list has been restricted since Android 11, which is just one year after the release of Android 10 (google2023privacy11, ). In other words, SDK’s privacy collection via the app list is infeasible on most devices nowadays. Nonetheless, given the number of tainted traces does not experience a significant change between the two assessments, SDK developers’ privacy compliance still seriously concerns us since their latest implementation still implies their persistent attempt to excessively and stealthily collect users’ privacy data.
8. Discussion
8.1. Causes of Privacy Mishandling
@@
An SDK’s access to privacy data is determined by permissions granted to its embedded apps. Third-party apps can obtain privileged permissions if pre-installed by OEMs. A malicious SDK embedded in a pre-installed app can exploit the system settings injection issues and execute privilege free-riding attacks, creating and sharing UUIs via the system settings. As reading system settings requires no permissions, privilege free-riding significantly undermines Google’s privacy protection efforts in restricting third-party app access to UUIs since Android 10.
@@
The release of Android apps faces strict regulation. Apps’ non-compliance in privacy preservation would result in removal from app stores or fines. However, SDK release lacks such fine-grained regulation and sufficient legislation to ensure compliance with personal data protection laws. The CAICT repository serves only as a publishing platform and search engine for available SDKs, without hosting privacy disclosures/policies. The Google Play SDK Index allows developers to voluntarily provide privacy policies but does not host binaries or enforce strict policies like its app store. Not all listed SDKs are “registered” in the Google Play SDK Console, meaning no guarantee of compliance with basic terms and conditions. Notably, even some Google in-house SDKs lack privacy policies on the SDK Index.
Furthermore, there is no platform-level compliance checking mechanism to ensure complete permission disclosure. The SDK Index only displays reminders about privacy-sensitive permissions, implying developers must undertake ultimate responsibility for privacy protection, regardless their code is self-implemented or imported from external parties.
8.2. Our Recommendations
@@
We regard that the pre-installed privileges open an attack surface for a malicious SDK, specifically when the SDK uses app-differential Android ID to identify a device user, transforms it into a de facto UUI, and provides free access with all apps installed on the victim device. Based on this finding, we advocate Google restricting pre-installed apps from writing data into the system settings in the future. In fact, pre-installed apps are free to create a customized UUI and share it even if it has not read any privacy data. Stop automatically granting WRITE_SETTINGS and instead allow apps with “preinstalled” privilege to explicitly request it at runtime can mitigate the risk of system settings injection. We also advise OEM manufacturers to minimize the number of pre-installed apps on their devices.
@@
We advocate transparent management over the permission requested by the embedding SDKs. Thus, app developers can have their own discretion in granting permissions without compromising functionality. We are delighted to see Google’s efforts in improving the SDK’s governance. A new mechanism called SDK Runtime has been introduced in Android 13 that provides a dedicated execution environment for SDKs (google2024sdk, ). However, this mechanism is not compulsory and still faces many usability and functionality limitations at the moment of this work being performed. Besides that, how to ensure the cooperation of SDK developers remains a challenging issue, given many SDKs are released and maintained in non-Google platforms.
8.3. Limitations

### v2

@@ IV-B RQ2: Privacy Collection and Potential Leakages at the SDK Level
As shown in Table IV, location and sensors’ data (category C3) reserves the greatest portion of collection actions, contributing 1002 out of 2318 (43.2%) caught collection actions. The remaining privacy data categories ordered by their number of caught collection actions are data for wireless communication (C2, 22.8%), data for chip, cellular and peripheral (C1, 17.9%), media and software-specific data (C4, 13.4%), and personal data (C5, 2.8%), respectively.
From the privacy data sharing perspective, the access of wireless communication data (C2) constitutes the largest part of the tainted flow, as 172 out of 346 (49.7%) traces are tainted. The numbers of sharing actions involving data in categories C3 and C1 are following up, reserving 37.6% and 12.7% of all sharing actions, respectively. No sharing actions about data in categories C4 and C5 are observed. From the identifiability perspective, we find that most privacy data involved in sharing actions are referred to as personally identifiable information (PII) in relevant literature [25, 26, 47, 48].
We remark that whether SDKs can collect privacy data eventually depends on the permission and privilege of their embedding apps, and tainted traces do not automatically imply the exfiltration in runtime. A manual validation is performed to exclude false positive tainted traces. We will elaborate more on them in the remainder of this section.
@@
Figure 3: A false positive example of SSID reading (line 7)
@@
During our analysis, we noticed that static taint analysis cannot well determine the execution paths in conditional and branched code. If any tainted trace is found spanning over an unreachable, dead, or logically infeasible code, we exclude the tainted trace from the statistics of potential leakages. In Figure 3, we provide an example code reverse-engineered from one of our examined SDKs named “mipush”. We find it attempts to read the SSID of the connected WiFi but is later assessed by us as a false alarm. The SSID reading action (can be realized by the method getExtraInfo (line 7) until Android 10) would only take place if the device is not connecting to WiFi (after line 4), therefore we determine this is a logically infeasible code regarding privacy data access. In this study, we manually exclude the false alarms from the 346 traces and accordingly identify 338 (97.7%) logically valid traces, among which each trace indicates an occurrence of privacy leakage issue.
For the sharing operations, we observe that 332 out of 338 (98.2%) valid traces are tainted at network interfaces. The six remaining traces are tainted at system setting APIs. Compared with sharing privacy data through the Internet, we remark that writing privacy data in system settings is a more worrying issue because it exploits the privilege of pre-installed apps and is capable of sharing user privacy even without reading any privacy data from the Android OS. We will detail this in Section V.
IV-C RQ3: SDK-Level Privacy Compliance
We resort to cross-checking the privacy policies with the actual data collection behaviors of the collected SDKs to understand the status quo of privacy compliance at the SDK level. We remark our assessment only focuses on the 109 out of 158 (69.0%) SDKs that provide privacy policies, and the 27 data types that appear in the data collection behaviors observed from the taint analysis. We present our assessment results in Table V.
TABLE V: Assessment of privacy compliance by data types
@@ IV-C RQ3: SDK-Level Privacy Compliance
Our evaluation finds that 40 out of 109 (36.7%) SDKs collected excessive privacy data compared with their privacy policies (e.g., an SDK collects personal identifiers but does not declare them in its privacy policy). Among the 40 SDKs that excessively collect users’ privacy, the number of undeclared data types ranges from 1 to 9. Each SDK has an average of three undeclared data types. As shown in Table V, 22 out of the 41 pre-identified data types (81.5%) have been excessively collected by the examined SDKs. The app list, clipboard data, and location are the top three most excessively collected data types. Specifically, 22 SDKs (20.2%) collect the app list without any declaration in their privacy policies, followed by clipboard data and location that are excessively collected by 11 (10.0%) SDKs. By jointly considering data collection behaviors (refer to Table IV), our findings show that the app list and clipboard data are the two most stealthily collected data types. Only 4 SDKs collecting the app list have mentioned it in their privacy policies (leading to 22 excessive collection behaviors). None of the 11 caught SDKs have declared to read the clipboard data.
@@
Our study reveals two key findings. First, some SDK developers use vague or broad terms like “identifiers” or “device information” instead of specifying the exact data being collected, leading to differing interpretations. For instance, while most developers use “identifiers” to indicate device identifiers, some refer to advertisement identifiers, causing over-claims for device identifiers and under-claims for Android ID or Google Ad ID. Second, although many SDKs declare access to both GPS and IP addresses for location data, our taint analysis shows that most SDKs only use location APIs, leading to over-claiming cases of IP addresses. This lack of clarity can confuse users about what data is actually collected, reducing transparency in privacy policies.
IV-D Privacy Compliance Re-inspection
On the basis of our large-scale analysis performed on 158 SDKs collected in October 2022, we re-visited the two sources and collected the same group of SDKs in their latest release version as of October 2023 and conducted a cross-version analysis. Table VI depicts a summary of our re-inspection results.
@@ IV-D Privacy Compliance Re-inspection
TABLE VI: Changes in the tainted traces after 12 months (only tainted data types are displayed)
@@
We find 311 traces tainted from the latest version of 158 SDKs. Compared with our initial assessment, we observed 35 (10.1%) fewer traces that attempt to collect and share user privacy data in public spaces.
@@
The change between the outcomes of the two assessments mainly falls in the sources of tainted traces. As shown in Table VI, we find obvious growth from the sharing of data relevant to media and installed software (C4), reserving over 1/3 of the total caught traces. Among the 105 tainted traces, reading and sharing the list of installed apps ranks the top among all sources, with 53 traces caught. Although the access to data in that category has been pervasively observed in our initial assessment, our re-inspection did not catch any traces of sharing such data.
Additionally, we find that the collection behaviors for location and SSID/BSSID have significantly declined, marking the two greatest decreases. There is also an overall reduction in the collection of UUIs, including data types in categories C1 and C2. Upon reviewing the historical changes in permission requirements, we note that most data types showing decreased collection behaviors had their access rules significantly updated in Android 10 [21], which imposed strict restrictions on third-party apps’ access to personally identifiable information. In contrast, we find the growth of tainted traces for the app list. Although the access to app list has also been restricted in a future release, these results still seriously concern us since the latest version of SDKs implies their persistent attempt to excessively and stealthily collect user privacy data.
V System Settings Injection Issues
Figure 4: Information flow demonstrates how the device ID is accessed and saved to the system settings by one of our analyzed SDKs (the obfuscated code is obtained from reverse engineering and has been simplified to save space)
TABLE VII: System setting entries found that contains a UUI
SDK Entry name Description of the value
[…]
Baidu LBS com.baidu.deviceid.v2 Package name and Android ID, MD5 $\ddagger$
[…]
$\dagger$ depending on the permissions granted to the embedded app.
$\ddagger$ inferred description based on the reading logic in the implementation.
During our analysis, we find some traces from an SDK named “Alicloud push” are sunk at writing a value into the system settings. Specifically, two entries of system settings are involved, as shown in Table VII. Writing into the system settings requires privileged permission so that ordinary third-party apps are unable to do so. However, we find that SDKs being embedded by pre-installed apps can write arbitrary values into the system settings even without the user’s awareness. We name this finding a system settings injection issue, because it can be manipulated by an SDK to create and share user-unresettable identifiers (UUIs).
Although leaking privacy data through covert channels has been studied in existing literature [49], the system settings injection issues still seriously concern us as the caught SDK subtly evades accessing highly sensitive UUIs (e.g., IMEI, as discussed in  [49]) but attempts to generate UUIs on their own. The caught SDK then takes advantage of its popularity and abuses the “preinstalled” privilege by counting on itself to be embedded in pre-installed apps or even privileged system apps. By exploiting this issue, the system settings become a covert channel to enable privacy-sensitive data sharing to apps without privileged permission.
Considering the exploitability of the pre-installed privileges, we carry out another round of investigation of all 158 SDKs to find if there exists any other potential privacy infringement by taking advantage of the system settings mechanism. As a result, we managed to find another two entries that have been read in an SDK named “Baidu LBS” (see Table VII). As their names imply, these two entries are obviously written by third-party privileged apps. We suspect the formation of the two entries should follow a similar routine, and accordingly reverse engineer the involved SDK and present our inferred descriptions of the two entries in Table VII. Unfortunately, we did not find any app to write these entries into the system settings. Who exploited this system settings issue remains unknown to us.
Impact
We notice that the two studied SDKs are widely adopted in third-party apps, especially in the Chinese market. Therefore, we conduct a small-scale study to explore the impact of the system settings injection issues. We collect devices from 9 OEM manufacturers (8 OEM devices purchased in the Chinese market [^4] These devices are manufactured by Honor, Huawei, Lenovo, OnePlus, Oppo, Samsung, Vivo, and Xiaomi. and 1 Google Pixel device purchased in the US market), download the packages of pre-install apps, and investigate whether the two aforementioned SDKs have been once embedded in the pre-installed apps. We observe a worrisome fact that except for the Google Pixel devices, all the tested devices have at least one pre-installed app using the investigated SDKs. According to the latest market report [50], these 8 OEM manufacturers reserve over 66% global Android smartphone market share. They also shipped over 94% new Android devices in China in 2023 [51]. This implies that almost all Android smartphones sold in the Chinese market may already pose a privacy leakage risk based on an assumption that the choice of pre-installed apps should remain the same across different models produced by a manufacturer.
VI Discussion
VI-A Causes of Privacy Mishandling
@@
An SDK’s access to privacy data depends on the permissions granted to the apps it’s embedded in. However, third-party apps pre-installed by OEMs can obtain privileged permissions. A malicious SDK in such an app can exploit system settings to deploy privilege free-riding attacks, creating and sharing UUIs without users’ awareness. As reading system settings require no permissions, privilege free-riding significantly undermines Google’s efforts in restricting third-party app access to UUIs since Android 10.
@@
It is well known that the release of Android apps is subject to strict regulations, with non-compliance in privacy preservation potentially leading to removal from app stores or fines. However, SDK releases lack similar fine-grained regulations and sufficient legislation to ensure compliance with personal data protection laws. Ironically, even some Google in-house SDKs do not have privacy policies listed on the SDK Index [^5] An example can be found at https://play.google.com/sdks/details/com-google-android-gms-play-services-auth. Currently, both repositories function merely as search engines for available SDKs, without a platform-level compliance checking mechanism to ensure complete permission disclosure.
VI-B Our Recommendations
@@
We regard that the pre-installed privileges open an attack surface for a malicious SDK, specifically when the SDK uses app-differential Android ID to identify a device user, transforms it into a de facto UUI, and provides free access with all apps installed on the victim device. Based on this finding, we advocate Google restricting pre-installed apps from writing data into the system settings in the future. We also advise OEM manufacturers to minimize the number of pre-installed apps on their devices.
@@
We advocate transparent management over the permission requested by the embedding SDKs. Thus, app developers can have their own discretion in granting permissions without compromising functionality. We are delighted to see Google’s efforts in improving the SDK’s governance. A new mechanism called SDK Runtime has been introduced in Android 13 that provides a dedicated execution environment for SDKs [52]. However, this mechanism is not compulsory and still faces many usability and functionality limitations at the moment of this work being performed. Besides that, how to ensure the cooperation of SDK developers remains a challenging issue, given many SDKs are released and maintained in non-Google platforms.
VI-C Limitations